# Forecasting Fold Bifurcations through Physics-Informed Convolutional Neural Networks

### v1

@@ 5.3 Polar transformation, normalization and moving mean filter
Figure 8: Prediction diagrams obtained by the CNN trained with data normalized in polar form filtered with a moving mean. Color interpretation: green-far, red-close, blue-after.
@@ 5.4 Polar transformation, normalization and logarithmic scale
Figure 9: Prediction diagrams obtained by the CNN trained with data normalized in polar form and in logarithmic scale. Color interpretation: green-far, red-close, blue-after.
The transformation into a logarithmic scale reflects the observation that the decrement of the oscillation amplitude is logarithmic in a linear system. Accordingly, if far from the bifurcation the nonlinearities are weak, the amplitude is expected to decrease approximately logarithmically (of course, modal interaction modifies this behavior), which in a logarithmic scale appears linear. Training the network with data in polar coordinates, normalized and scaled in logarithmic coordinates provided better results than seen in the previous case for the nonlinearly damped oscillator (Fig. 8(a)), which were excellent for all the investigated range of $c_{1}$. However, the network provided a completely wrong classification for the mass-on-moving-belt system (Fig. 8(b)); for this system, all far cases were incorrectly classified as after. Interestingly, some close trajectories were correctly classified. It is hard to explain understand the mechanism leading to such a wrong classification. Regarding the van der Pol-Duffing oscillator (Fig. 8(c)), the network performed well; in this case, the region of close trajectories was much larger than for the networks whose data were not in logarithmic scale (Figs. 6(c) and 7(c)); the prediction based on the primary system displacement ($x_{1}$) provided a smaller close region than if based on the vibration absorber displacement ($x_{2}$). About the pitch-and-plunge wing profile, the classification is acceptable; however, some trajectories in the close region were mistakenly classified as after; this problem is less significant if classification is based on plunge motion ($y$), rather than pitch ($\alpha$).
@@ 5.5 Polar transformation, normalization, logarithmic scale and moving mean
Figure 10: Prediction diagrams obtained by the CNN trained with data normalized in polar form and logarithmic scale and then filtered with a moving mean. Color interpretation: green-far, red-close, blue-after.
@@
The last case we consider is similar to the previous one, with an additional moving mean filter applied to the data to remove variations of the amplitude in polar coordinates during each complete oscillation. The same filter was applied for the case not in logarithmic scale and produced no improvement (see Sect. 5.3). Conversely, in this case, it significantly enhanced the accuracy of the network. All tested cases provided very good results. In particular, trajectories of the nonlinearly damped oscillator were correctly classified for all the investigated range of $c_{1}$ (Fig. 9(a)). Excellent classification was also provided for the mass-on-moving-belt system (Fig. 9(b)); we remark that this was the only normalization that enabled the network to correctly classify close trajectories for this system. The network was able to correctly distinguish between the three classes also for the van der Pol-Duffing oscillator (Fig. 9(c)) and pitch-and-plunge wing (Fig. 9(d)), using any of the coordinates.
@@ 6 Discussion and conclusions
The performed analysis demonstrates that a properly trained neural network can recognize a trajectory passing close to a fold bifurcation. However, in order for the network to capture the important features of such trajectories and provide correct classification for different systems, it is essential that training data are properly normalized. This step eliminates inessential information from the data, such as oscillation frequency, allowing the network to focus on important ones, like the decrement of oscillation amplitude. The results obtained with this method are remarkable. In fact, a network trained with a single, very simple system performed very well on three different mathematical systems modeling mechanical systems of engineering relevance, thus exhibiting excellent extrapolation capabilities. The most important step of the normalization was the transformation in polar coordinates, which eliminates almost completely information about oscillation frequency, thus highlighting amplitude. Applying a logarithmic scaling, providing a physically more relevant scale, further improved the network performance. Finally, introducing a moving mean filter and eliminating any residual information about variations in single oscillation periods enhanced even more correct classification capabilities. The moving mean could also eliminate noise in the data, an aspect not investigated in the present study. Every step of data normalization is related to the physical understanding of the problem, and it is not standard for signal processing, except for the moving mean.
@@
We note that there is a significant overlap in the region where trajectories are classified as either far or close (particularly for the van der Pol-Duffing oscillator). This is mostly because some initial conditions lead to a dynamical behavior almost unaffected by the fold, while other initial conditions generate trajectories more distorted by it. This phenomenon was clearly illustrated in habib2023predicting .
Regarding the limitations of the approach, we remark that the range of the initial conditions significantly affects the network’s accuracy. In fact, referring to the best-performing case, the network has to distinguish between scenarios similar to those in Figs. 4(j) and 4(k); Fig. 4(k) might be distorted if large initial conditions move down the bump, confusing the CNN.
Additionally, we note that in the studied cases, the simulations were run until the systems reached extremely small oscillation amplitudes (about $e^{-10}$, see Fig. 4(j)). In the case of a real system, and in the presence of noise, it is unreasonable to expect the system to reach such small amplitudes, which would modify the numerical values in the far and close cases, potentially generating wrong classifications. In particular, far and close trajectories might be classified as after.
These potential limitations can be overcome by increasing the heterogeneity of training data, for example, by adding noise and considering different systems. Another way to solve the issues is to use a sliding window convolutional neural network, which would analyze the signal by dividing it in time window frames. This way, the network would provide a list of outputs for each signal. This output list could then be further analyzed to obtain a single classification output.
@@
At last, we note that this study is limited in its scope, as the network only distinguishes between trajectories far, close, or after the fold and does not try to estimate the position of the fold, as other deterministic methods do lim2011forecasting ; habib2023predicting . While training the network to identify the position of the fold based on some time series computed before the fold is probably very challenging, it is likely feasible to add more classes to distinguish, for example, between close and very close. Additionally, including trajectories starting at lower amplitude, it might be possible to distinguish also between trajectories far from, close to, or after the Andronov-Hopf bifurcation (if present). However, these further developments probably require more diverse and abundant training data.
In general, we think that this study paves the way for the development of an algorithm for fold bifurcation prediction applicable in real-life problems, which might represent a useful tool in engineering, especially for safety monitoring of dynamical systems.

### v2

@@ 5.3 Polar transformation, normalization and moving mean filter
Figure 9: Prediction diagrams obtained by the CNN trained with data normalized in polar form filtered with a moving mean. Each point corresponds to a different time series. Dashed lines indicate the fold bifurcation. The colors mark the predicted classification. Color interpretation: green-far, red-close, blue-after.
@@ 5.4 Polar transformation, normalization and logarithmic scale
Figure 10: Prediction diagrams obtained by the CNN trained with data normalized in polar form and in logarithmic scale. Each point corresponds to a different time series. Dashed lines indicate the fold bifurcation. The colors mark the predicted classification. Color interpretation: green-far, red-close, blue-after.
The transformation into a logarithmic scale reflects the observation that the decrement of the oscillation amplitude is logarithmic in a linear system. Accordingly, if far from the bifurcation the nonlinearities are weak, the amplitude is expected to decrease approximately logarithmically (of course, modal interaction modifies this behavior), which in a logarithmic scale appears linear. Training the network with data in polar coordinates, normalized and scaled in logarithmic coordinates provided better results than seen in the previous case for the nonlinearly damped oscillator (Fig. 10(a)), which were excellent for all the investigated range of $c_{1}$. However, the network provided a completely wrong classification for the mass-on-moving-belt system (Fig. 10(b)); for this system, all far cases were incorrectly classified as after. Interestingly, some close trajectories were correctly classified. It is hard to understand the mechanism leading to such a wrong classification. Regarding the van der Pol-Duffing oscillator (Fig. 10(c)), the network performed well; in this case, the region of close trajectories was much larger than for the networks whose data were not in logarithmic scale (Figs. 8(c) and 9(c)); the prediction based on the primary system displacement ($x_{1}$) provided a smaller close region than if based on the vibration absorber displacement ($x_{2}$). However, the overlapping region is much larger than in the previous cases (Table 2). About the pitch-and-plunge wing profile, the classification is acceptable; however, some time series in the close region were mistakenly classified as after; this problem is less significant if classification is based on plunge motion ($y$), rather than pitch ($\alpha$).
@@ 5.5 Polar transformation, normalization, logarithmic scale and moving mean
Figure 11: Prediction diagrams obtained by the CNN trained with data normalized in polar form and in logarithmic scale and then filtered with a moving mean. Each point corresponds to a different time series. Dashed lines indicate the fold bifurcation. The colors mark the predicted classification. Color interpretation: green-far, red-close, blue-after.
@@
The last case we considered is similar to the previous one, with an additional moving mean filter applied to the data to remove variations of the amplitude in polar coordinates during each complete oscillation. The same filter was applied for the case not in logarithmic scale and produced no improvement (see Sect. 5.3). Conversely, in this case, it significantly enhanced the accuracy of the network. All tested cases provided very good results. In particular, trajectories of the nonlinearly damped oscillator were correctly classified for all the investigated range of $c_{1}$ (Fig. 11(a)). Excellent classification was also provided for the mass-on-moving-belt system (Fig. 11(b)); we remark that this was the only normalization that enabled the network to correctly classify close trajectories for this system. The network was able to correctly distinguish between the three classes also for the van der Pol-Duffing oscillator (Fig. 11(c)) and pitch-and-plunge wing (Fig. 11(d)), using any of the coordinates. In terms of limitations, we note that the overlapping region is relatively large in all cases, especially for the pitch-and-plunge system (Table 2).
6 Results with noisy signals
(a) Original/noisy signals
(b) Polar-log transf. ($u=0.902$)
(c) Polar-log transf. ($u=0.909$)
Figure 12: Comparison of the original time series obtained from direct numerical integration (red) and signal with added Gaussian noise (black) ($\text{SNR}=15$). (a) Signal from the pitch-and-plunge wing profile ($u=0.9024$); (b) signals in (a) after transformation into polar coordinates, normalization, and transformation into logarithmic scale; (c) signals as in (b) but with $u=0.90905$.
In this section, we investigate the performance of the trained network with noisy data. The analysis was limited to the most successful signal pre-processing strategy according to the analysis performed so far, namely polar transformation, normalization, transformation into natural logarithmic scale, and moving mean filtering (however, we note its poor performance in terms of overlapping region (Tab. 2)). The same data utilized in the previous analysis was corrupted with a Gaussian noise using a signal-to-noise-ratio (SNR) of 15 dB. The corrupted signals were then post-processed as described in Sect. 4.2. We note that the post-processing combines position and velocity signals. For this analysis, the noise was added independently to both position and velocity signals, which might generate artifact results; however, this is not expected to compromise the qualitative results of the analysis. Noisy signals were obtained through the MATLAB built-in function awgn(x,SNR,‘measured’), which automatically measures the signal average power and, accordingly, defines the power of the added noise. The signal power was measured independently for each signal, leading to higher absolute noise levels for signals with higher power.
A comparison of a signal with and without noise is depicted in Fig. 12(a). The depicted signal refers to the pitch-and-plunge wing profile with $u=0.9024$, which is close to the fold bifurcation ($u^{*}=0.911$). Despite the moderate SNR, it is observed that the original signal is practically unrecognizable for low-amplitude oscillations.
Figure 12(b) compares the clean and noisy signals provided to the network. The two signals are relatively close to each other at large amplitude, while the logarithmic scale enhances the difference between them at low amplitude.
(a) Nonlinear damping system
(b) Mass-on-moving-belt
(c) Van der Pol-Duffing oscillator
(d) Pitch-and-plunge wing
Figure 13: Prediction diagrams obtained by the CNN trained with data normalized in polar form and logarithmic scale and then filtered with a moving mean. A Gaussian noise with an SNR of 15 was added to the test data (not the training data). Each point corresponds to a different time series. Dashed lines indicate the fold bifurcation. The colors mark the predicted classification. Color interpretation: green-far, red-close, blue-after.
For time series of the same system and in the same parameter range of the training data (the oscillator with nonlinear damping with $c_{1}=0.5$), but with added noise, the accuracy of the NN decreased to 72.3 % (it was 100 % in the absence of noise). Testing the network for the same system, but with $c_{1}=0.1$, provided an accuracy of 71.5 % (also 100 % without noise). In both cases, all errors are relative to close trajectories erroneously classified as far.
Results of the analysis for the other systems are provided in Fig. 13. As a general trend, we note that the range of solutions classified as close is significantly reduced. This phenomenon occurs for all studied systems. In particular, no solution is classified as close for the mass-on-moving-belt system. For all systems, trajectories that are no longer classified as close are now classified as far. The overlapping between far and close regions is also very large (Table 2).
Except for the mass-on-moving-belt system, this result cannot be strictly considered erroneous because of the arbitrary distinction between far and close cases. Nevertheless, from a practical perspective, it is unsatisfactory because it makes the prediction non-conservative concerning the fold bifurcation identification.
Because of the black-box nature of the method, it is not trivial to understand why the NN reduces the range of close solutions in the presence of noise. Looking at Figs. 5(j) and 5(k), it can be recognized that the NN distinguishes far and close trajectories from the bump present in close trajectories. Figure 12(b) depicts a trajectory classified either as close or far in the absence or presence of noise, respectively. Conversely, Fig. 12(c) shows a trajectory closer to the fold, classified as close in both cases. We hypothesize that the noise reduces the slope of the signal, making the bump less visible. However, for trajectories sufficiently close to the fold, as the one in Fig. 12(c), the bump is pronounced enough to be recognized by the NN even in the presence of noise, which classifies the trajectory as close. Nevertheless, it is hard to validate this hypothesis. Additionally, we note that, since the noise level is proportional to the signal level, close trajectories present a relatively high noise level because their power level is initially high, making noise significant for the second part of the signal when the vibration amplitude decreases. This effect might contribute to the wrong classification of close trajectories.
[…]
The NN trained with noisy data provided a 100 % accuracy for data in the same parameter range as the training data, confirming the 100 % accuracy also for data from the same system, but with $c_{1}=0.1$. This result is particularly surprising and already represents a significant improvement compared to the NN trained with clean data. It implies that, despite the presence of noise, the time series from the three categories are still different enough for the NN to correctly classify them with no error. Also in this case, if the gap between far and close categories was smaller, the accuracy would probably decrease. We note that the time series utilized for the validation are obtained for exactly the same parameter values as for the case without noise. The specific $c_{3}$ values utilized are illustrated in Fig. 6
(a) Nonlinear damping system
(b) Mass-on-moving-belt
(c) Van der Pol-Duffing oscillator
(d) Pitch-and-plunge wing
Figure 14: Prediction diagrams obtained by the CNN trained with data normalized in polar form and logarithmic scale and then filtered with a moving mean. A Gaussian noise with an SNR of 15 was added to the training and test data. Each point corresponds to a different time series. Dashed lines indicate the fold bifurcation. The colors mark the predicted classification. Color interpretation: green-far, red-close, blue-after.
Figure 14 illustrates the predictions obtained by the NN for trajectories regarding the other systems. For the oscillator with nonlinear damping (Fig. 14(a)), results are satisfactory for the whole range of parameter values investigated. Also regarding the mass-on-moving-belt system (Fig. 14(b)), the NN was able to identify trajectories in all classes, with some overlapping between close and far trajectories. The same occurs for the pitch-and-plunge airfoil (Fig. 14(d)), with data from any coordinates. For this system, the overlapping region is particularly small (Table 2), which marks an improvement also with respect to the analysis with clean data. The only system for which the NN did not work well is the van der Pol-Duffing oscillator, where predictions obtained from $x_{1}$ confused almost all far trajectories as close. Interestingly, this error is opposite to the one done by the NN trained with clean data.
At last, the performance of the NN trained with noisy data was tested on trajectories without noise. The NN was able to correctly predict the classification for all the systems, including both coordinates of the van der Pol-Duffing oscillator. The relative figure is omitted for the sake of brevity.
Considering the limited data utilized for the training and the absence of specific measures for filtering the noise in the data, we believe that the results are encouraging concerning the effectiveness of the proposed methodology.
@@ 7 Discussion and conclusions
The performed analysis demonstrates that a properly trained neural network can recognize a trajectory passing close to a fold bifurcation. However, in order for the network to capture the important features of such trajectories and provide correct classification for different systems, it is essential that training data are properly normalized. This step eliminates inessential information from the data, such as oscillation frequency, allowing the network to focus on important ones, like the decrement of oscillation amplitude. The results obtained with this method are remarkable. In fact, a network trained with a single, very simple system performed very well on three different mathematical systems modeling mechanical systems of engineering relevance, thus exhibiting excellent extrapolation capabilities. The most important step of the normalization was the transformation in polar coordinates, which eliminates almost completely information about oscillation frequency, thus highlighting amplitude. Applying a logarithmic scaling, providing a physically more relevant scale, further improved the network performance. Finally, introducing a moving mean filter and eliminating any residual information about variations in single oscillation periods enhanced even more correct classification capabilities. Every step of data normalization is related to the physical understanding of the problem, and it is not standard for signal processing, except for the moving mean.
@@
We observe a significant overlap in the regions where trajectories are classified as either far or close. This overlap arises primarily because some initial conditions result in dynamical behavior that is minimally affected by the fold, while others produce trajectories that are more significantly distorted by it. This phenomenon was clearly demonstrated in [45]. Furthermore, we note that the overlapping region is larger for the best-performing data pre-processing strategy. In contrast, pre-processing that involves only transformation into polar form, without applying a logarithmic scale, results in moderate overlap. This observation suggests that the benefit of the logarithmic transformation cannot be conclusively established. Moreover, this might reflect an inherent limitation of the approach, which requires a compromise between achieving clear boundaries at the cost of higher misclassification rates, or tolerating fuzzier boundaries with improved accuracy. To determine the optimal pre-processing strategy and to minimize fuzziness, more comprehensive testing is needed, including training on a diversified set of systems, and the analysis of various NN architectures.
Regarding the limitations of the approach, we remark that the range of the initial conditions significantly affects the network’s accuracy. In fact, referring to the best-performing case, the network has to distinguish between scenarios similar to those in Figs. 5(j) and 5(k); Fig. 5(k) might be distorted if large initial conditions move down the bump, confusing the CNN.
These potential limitations can be overcome by increasing the heterogeneity of training data, for example, by considering different systems. Another way to solve the issues is to use a sliding window convolutional neural network, which would analyze the signal by dividing it in time window frames. This way, the network would provide a list of outputs for each signal. This output list could then be further analyzed to obtain a single classification output.
In terms of possible applications, this approach is implementable only on systems that can be forced to undergo, or naturally experience, large perturbations, such as, for example, an adaptive cruise control car-following system experiencing a sudden cut-in [88]. Although this limitation is common for most methods available for fold bifurcation identification [42, 45], it still significantly limits the range of applications. This aspect should be addressed in future studies. Besides, the bifurcation parameter is assumed quasi-static, i.e., constant for each trajectory. Trajectories with varying bifurcation parameters might include additional challenges.
Another limitation worth mentioning is that, in the bistable region, only trajectories approaching the periodic solutions were considered. Future developments of this study will include the analysis of trajectories converging to the equilibrium solution in the bistable region.
The trained NN exhibited a sensitivity to noise in the signals, which resulted in a degradation of accuracy; however, this sensitivity was significantly mitigated by including noise also in the training data.
@@
Regarding the implementation of the method to other systems, a clear limitation is in the bifurcation diagram: the method can be applied only to systems having a fold bifurcation with properties similar to those of the systems discussed here. Besides, since the division between close and far trajectories is arbitrary, special care should be taken while applying the method to a new system.
At last, we note that this study is limited in its scope, as the network only distinguishes between trajectories far, close, or after the fold and does not try to estimate the position of the fold, as other deterministic methods do [42, 45]. While training the network to identify the position of the fold based on some time series computed before the fold is probably very challenging, it is likely feasible to add more classes to distinguish, for example, between close and very close trajectories. Additionally, including trajectories starting at lower amplitude, it might be possible to distinguish also between trajectories far from, close to, or after the Andronov-Hopf bifurcation (if present). However, these further developments probably require more diverse and abundant training data.
In general, we believe that this study advances the development of an algorithm for fold bifurcation prediction applicable to real-world problems, which could serve as a valuable tool in engineering, particularly for safety monitoring of dynamical systems. However, further research is needed to address the limitations discussed and to make the method practically implementable.